What is the standard equation and state whether it is an ellipse, parabola, circle, or hyperbola x^2+y^2+6x=7
Answer by stanbon(48569)

What is the standard equation and state whether it is an ellipse, parabola, circle, or hyperbola x^2+y^2+6x=7
------------------
Complete the square separately for the x and the y terms, as follows:
x^2+6x+9 + y^2 = 7+9
Factor as follows:
(x+3)^2 + (y-0)^2 = 16
-----------
this is a circle with center at (-3,0) and radius = 4
===========
Cheers,
Stan H.